OpenStudy (anonymous):

Estimate the sum or difference. 8 12/31 + 4 5/6 A 13 B 13 1/2 C 14

TheSmartOne (thesmartone):

8 12/31= 8.387 which rounds to 8 4 5/6= 4.8333 which rounds to 5 So 8+5=?

OpenStudy (anonymous):

13

TheSmartOne (thesmartone):

Exactly